OpenStudy (anonymous):

Write the equation of a line in slope-intercept form that has a slope of -2 and passes through the point (2, -8).

OpenStudy (marissalovescats):

You are going to plug the point and your slope into the equation: y-y1=m(x-x1) Then from there you want to put it into standard form y=mx+b

OpenStudy (marissalovescats):